OpenStudy (luigi0210):

Integration, I showed what I did, just need someone to verify that it's right:

OpenStudy (luigi0210):

\[\LARGE \int\frac{dx}{cosx-sinx+1}\] \(\Large z=tan\frac{x}{2}\) so \(\Large dx=\frac{2dz}{1+z^2}\) Sub in: \[\LARGE \int\frac{\frac{2dz}{1+z^2+1}}{\frac{1-z^2}{1+z^2}-\frac{2z}{1+z^2}+1}\] Multiply by \(1+z^2\) \[\LARGE 2\int \frac{dz}{1-z^2-2z+1+z^2}\] Simplify it \[\LARGE 2\int \frac{dz}{2-2z}\] factor and simplfy \[\LARGE \int \frac{dz}{1-z}\] \[\LARGE -ln|1-z|+C\] \[\LARGE -ln|1-tan\frac{x}{2}|+C\]

geerky42 (geerky42):

You lost me at "Luigi0210"

OpenStudy (anonymous):

You lost me at Hi

OpenStudy (anonymous):

Fine, you lost me at ∫

OpenStudy (anonymous):

u=tan(x/2), du = 1/2 sec^2(x/2)dx \[sinx = \frac{ 2u }{ u^2+1 }\] \[cosx = \frac{ 1-u^2 }{ u^2+1 }\] \[dx = \frac{ 2du }{ u^2+1 }\] \[\huge \int\limits \frac{ 2 }{ (u^2+1)(\frac{ 2u }{ -u^2+1 }+\frac{ 1-u^2 }{ u^2+1 }+1) }du\]

geerky42 (geerky42):

Once again, we are saved by Batman!

OpenStudy (anonymous):

\[ = \int\limits \frac{ 1 }{ 1-u }du, ~~~ t = 1-u, dt = -du \implies -\int\limits \frac{ 1 }{ t }dt\]

OpenStudy (anonymous):

\[-\ln(1-u)+c \implies -\ln(1-\tan \frac{ x }{ 2 })+C\] Good job, but there is a better answer if you have restricted x values, not sure if you care for it though lol.